# On the Invariant and Geometric Structure of the Holomorphic Unified Field Theory

### v1

@@ 3 Real Slice Geometry
We will now explain real-slice geometry as a finer bundle, to show how the real slice geometry encodes symmetries in HUFT. By a finer bundle over $M$ we mean an enlargement of a given bundle $q:P\to M$ by adjoining an extra fiber $E_{y}\to M$, encoding the imaginary directions $y$ from the complexification $M_{\mathbb{C}}$:
@@
so that fiberwise $\widehat{P}_{x}\cong P_{x}\times(E_{y})_{x}$ and $\pi_{1}$ simply forgets $y$. In our setting $P=P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$, so $\widehat{P}=(P_{\mathrm{Spin}}\times_{M}P_{G})\times_{M}E_{y}$. This is not a reduction of the structure group nor a refinement of an atlas, but it is an enrichment of the fiber used to keep track of the $y$-directions while all fields and dynamics are evaluated on the real slice $M$. We let $M$ be a smooth oriented, time–orientable spin 44–manifold, the real slice. We have $M_{\mathbb{C}}$ as a complexification with coordinates $z^{\mu}=x^{\mu}+iy^{\mu}$ and projection $\pi:M_{\mathbb{C}}\to M$, $\pi(x,y)=x$. The fibers $Y_{x}:=\pi^{-1}(x)$ are real 44–planes isomorphic to $\mathbb{R}^{4}$, defining a rank–$4$ real vector bundle $E_{y}\to M$ with total space canonically identified near $y=0$ with a neighborhood of $M\subset M_{\mathbb{C}}$.
We have $P_{Spin}\to M$ as the $Spin(1,3)$ frame bundle and $P_{G}\to M$ a principal GG–bundle with $G=SU(3)\times SU(2)\times U(1)$. Now we define the product bundle $P_{tot}:=P_{Spin}\times_{M}P_{G}$ with structure group $H:=Spin(1,3)\times G$. We refine this by adjoining the yy–fiber:
@@
whose points encode local Lorentz frames, internal gauge frames, and the imaginary directions $y$ of the complexified manifold. Now we choose an HH–connection $\mathcal{A}=(\omega,A)$ on $P_{\mathrm{tot}}$ and an Ehresmann connection on $\pi:M_{\mathbb{C}}\to M$ that splits $TM_{\mathbb{C}}\simeq\mathsf{Hor}\oplus\mathsf{Ver}$ near $y=0$. Now we write the unified curvature:
@@
and let $h\in\Gamma(\text{Sym}^{2}T^{*}M)$ be a Lorentzian metric, used for index operations and Hodge duals. On $M_{\mathbb{C}}$ we consider a Hermitian field
@@
and identify on the real slice $B=F$ either by definition or via a holomorphic compatibility term enforcing BF=0B-F=0 in the equations of motion. The base metric $h$ transforms under $\text{Diff}(M)$ as its Levi–Civita and spin connection $\omega$ is the $Spin(1,3)$ part of $A$. The curvature $R$ is the gravitational field strength; all contractions, Hodge duals, and stress–energy are built using $h$. The internal symmetry $G=SU(3)\times SU(2)\times U(1)$ acts on $P_{G}$; the unified gauge potential $A=A^{(3)}\oplus A^{(2)}\oplus A^{(1)}$ and field strength $F=F^{(3)}\oplus F^{(2)}\oplus F^{(1)}$ live in $\Omega^{2}(M,\mathrm{ad}P_{G})$. On the real slice, the antisymmetric part of the Hermitian metric equals this curvature: $g_{[\mu\nu]}\equiv B_{\mu\nu}=F_{\mu\nu}$. Finer bundle with $y$ inside, The yy–fiber $E_{y}$ records the imaginary directions of $M_{\mathbb{C}}$. The Ehresmann splitting ties variations in $y$ to horizontal transport on $M$. In the holomorphic description, the mixed (x,y)(x,y)–geometry packages the gauge sector into the antisymmetric two–form $B$ while preserving standard spacetime geometry in $h$. By Noether I & II the single invariant action: A single $\text{Diff}(M)\times G$–invariant action:
@@
and by Noether’s second theorem the unified Bianchi identities $D_{\omega}R=0$, $D_{A}F=0$, together with covariant current conservation $\nabla_{\mu}T^{\mu\nu}=0$, $D_{A}J=0$. The same symmetry explains the conservation laws. Writing $G=SU(3)\times SU(2)\times U(1)$ with Lie algebra $\mathfrak{g}=\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$ and Killing form $\kappa$, we decompose:
@@
Figure 2: The upper ribbon is $M_{\mathbb{C}}$ with coordinates $z^{\mu}=x^{\mu}+i\,y^{\mu}$, the lower ribbon is the real slice $M$. The projection $\pi:M_{\mathbb{C}}\!\to M$ sends $(x,y)\mapsto x$. The black dots mark the point $s_{0}(x)=(x,0)\in M_{\mathbb{C}}$ (label $s$) and its projection $x\in M$ (label $\pi$), so that $\pi\circ s_{0}=\mathrm{id}_{M}$. The label $\mathbb{P}_{y}$ denotes the rank–$4$ real vector bundle $E_{y}\to M$ of imaginary directions: for each $x\in M$, the fiber $Y_{x}=\pi^{-1}(x)\cong\mathbb{R}^{4}$ carries coordinates $y^{\mu}$.
The diagram shows the complexification $\pi:M_{\mathbb{C}}\!\to M$, with the upper ribbon $M_{\mathbb{C}}$ and the lower ribbon the real slice $M$. The point labelled $s$ is the image of the zero section $s_{0}:M\hookrightarrow M_{\mathbb{C}}$, $s_{0}(x)=(x,0)$; the lower point labelled $\pi$ is the basepoint $x\in M$, so $\pi\!\circ s_{0}=\mathrm{id}_{M}$. The curved line between them indicates the fiber $Y_{x}:=\pi^{-1}(x)$ over $x$, which is a real 44–plane isomorphic to $\mathbb{R}^{4}$ and coordinatized by the imaginary directions $y^{\mu}$. We denote by $\mathbb{P}_{y}$ in the figure our $E_{y}$ in the text the rank–$4$ real vector bundle $E_{y}\to M$ whose fiber at $x$ is $Y_{x}$; near $y=0$ its total space is canonically identified with a neighborhood of $M\subset M_{\mathbb{C}}$. This bookkeeping of the yy–directions, lets us speak of a refined finer bundle over $M$ that remembers the imaginary displacements while all fields are ultimately evaluated on the real slice.
@@ 4 Symmetry Completeness and its Physical Equivalence
As a sanity check, we will show that we recover symmetry completeness and explore its physical equivalence. We have $M$ as an oriented, time-orientable spin 44–manifold and let $P_{\text{Spin}}\to M$ be the $\text{Spin}(1,3)$ frame bundle and $P_{G}\to M$ a principal GG–bundle with $G=SU(3)\times SU(2)\times U(1)$. Set $P_{\text{tot}}=P_{\text{Spin}}\times_{M}P_{G}$ with structure group $H=\text{Spin}(1,3)\times G$. Now let $h\in\Gamma(\text{Sym}^{2}T^{*}M)$ be a Lorentzian metric and $\mathcal{A}=(\omega,A)$ an HH–connection with curvature $\mathcal{F}=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=(R,F)$, where $R$ is the Riemann curvature $2$-form and $F$ the Yang–Mills field strength. On the complexified bundle of $(0,2)$-tensors, define the Hermitian field as in Eq( 53) and on the real slice impose the compatibility $B=F$ either as a definition or via a holomorphic penalty term enforcing BF=0B-F=0 in the equations of motion.
We consider the single $\text{Diff}(M)\times G$–invariant action:
@@
where $\langle\cdot,\cdot\rangle_{h}$ is the fiberwise inner product induced by $h$ and the Killing form on $\mathfrak{g}$, and $\Psi$ denotes matter fields. Here $\mathcal{A}=(\omega,A)$ with curvature $\mathcal{F}=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=(R,F)$, so $D_{\mathcal{A}}=(D_{\omega},D_{A})$ on $\mathfrak{spin}(1,3)\oplus\mathfrak{g}$.
Theorem 1
@@
    The Euler–Lagrange equations of $S$ are:
@@
    (Classical equivalence of physics) The map:
@@
    (Quantum equivalence, formal) If one includes a gauge-invariant penalty $\frac{\lambda}{2}\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$ and integrates out $B$, then, for $\lambda\to\infty$, the generating functional reduces to that of the standard EYM++matter theory, so perturbative correlators and SS–matrix elements agree.
Under HH–gauge transformations $F\mapsto g^{-1}Fg$ and the Killing form is Ad–invariant, so $\mathrm{tr}(F\wedge*_{h}F)$ is gauge invariant. Each term in (59) is a scalar density, hence $\text{Diff}(M)$–invariant. For a connection $A$ on a principal HH–bundle, $D_{A}F=0$ is the unified Bianchi identity, which splits as $D_{\omega}R=0$ and $D_{A}F=0$. Diffeomorphism invariance implies the contracted Bianchi identity $\nabla_{\mu}G^{\mu\nu}=0$ and hence $\nabla_{\mu}T^{\mu\nu}=0$ on shell GG–invariance implies $D_{A}J=0$. Varying $A$ with $h$ fixed gives:
@@
Given any EYM++matter solution $(h,A,\Psi)$, define $(g,A,\Psi)$ by $g=h+iF$ and $\mathcal{A}=(\omega(h),A)$. Then $(g,A,\Psi)$ solves the unified Euler–Lagrange system on the real slice because (60) hold, and $B=F$ by construction. Any unified real-slice solution $(g,A,\Psi)$ has $g=h+iB$ with $B=F$ from the BB–equation, and its $(h,A,\Psi)$ satisfies (60). Both constructions are natural with respect to $\text{Diff}(M)\times G$, so they descend to a bijection of quotient solution spaces. Classical observables built from $(h,A,\Psi)$, such as fluxes, charges, stress-energy, coincide with those built from $(g,A,\Psi)$ restricted to the real slice.
We add $S_{\lambda}[B,A;h]=\tfrac{\lambda}{2}\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$. The BB–integral is Gaussian:
@@
The consequences of this are that we have the same symmetries; $\text{Diff}(M)\times G$ invariance, Bianchi identities, and covariant current and stress-energy conservation are identical to GR++YM. The same observables, such as charges, fluxes, and classical predictions, match, so perturbative quantum correlators agree when $B=F$ is enforced. The unified packaging; gravity $R$ and gauge $F$ are components of one curvature $\mathcal{F}$ of one connection $\mathcal{A}$ on one bundle $P_{\mathrm{tot}}$.
Theorem 2
Let $(M,h)$ be a real Lorentzian 44–manifold admitting a complexification $M_{\mathbb{C}}$. Let $P_{\mathrm{tot}}\to M$ be a principal HH–bundle with $H=\mathrm{Spin}(1,3)\times G$ where $G$ is compact, reductive, and contains the SM gauge group. Consider fields $\big(h_{\mu\nu},\mathcal{A}=(\omega,A_{\mathrm{YM}}),\Psi,\Phi,g_{\mu\nu}\big)$ with:
@@
Assume the dynamics are invariant under the automorphism group:
@@
implemented off–shell by a nilpotent BRST differential $s$ (Noether II setting). The Lagrangian density is a diffeo and gauge–natural polynomial in the fields and their covariant derivatives, and may depend on the covariant d’Alembertian $D^{2}$ only through an entire functional calculus $F(D^{2}/M_{*}^{2})$ that commutes with $s$ and all covariant derivatives. In the limit $M_{*}\to\infty$, the quadratic operators for $(h,A,\Psi,\Phi)$ reduce to second order in derivatives, with indices raised and lowered solely by $h$. Reflection positivity on the Euclidean slice as no extra propagating poles other than those of $(h,A,\Psi,\Phi)$. No additional massless higher–spin gauge fields beyond spin–2 $h$ and spin–1 connections; interactions marginal or relevant in the IR. The local BRST anomaly cohomology $H^{1,4}(s\mid d)$ vanishes for the chosen matter representation, no gauge or diffeo anomalies.
Then, modulo ss–exact terms, total derivatives, and higher–dimension operators suppressed by $M_{*}$, the most general action is, after field redefinitions, equivalent to:
@@
with the following consequences; BRST covariance implies the antisymmetric piece $B_{\mu\nu}$ transforms as a curvature two–form, hence on the real slice:
$B_{\mu\nu}\;\equiv\;(F_{\mathrm{YM}})_{\mu\nu}\quad\text{(in an appropriate adjoint/trace)}\,$, (69)
so $g_{\mu\nu}=h_{\mu\nu}+iB_{\mu\nu}$ is a kinematic packaging and does not introduce an independent 2-form gauge sector. Preserving the full local symmetry off–shell to all loops restricts UV softening to entire functions $F(D^{2}/M_{*}^{2})$ at covariant kinetic operators and consistently at vertices. This renders perturbation theory UV–finite while the IR ($M_{*}\to\infty$) recovers local EYM++matter. Noether II yields the unified Bianchi and Slavnov–Taylor identities, and:
@@
with $T^{\mu\nu}$ the Belinfante tensor built from $h$. Up to $S_{\mathrm{top}}$, Euler, Pontryagin, $\theta$–terms, no further diffeo or gauge–natural, unitary, second–order couplings exist.
In this section, we specify the geometric data $P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$ with structure group $H=\mathrm{Spin}(1,3)\times G$, a single HH–connection $\mathcal{A}=(\omega,A)$ with unified curvature $\mathcal{F}=(R,F)$, and a Lorentzian metric $h$. The master action Eq. (7) is written and shown to be $\mathrm{Diff}(M)\times H$–invariant by construction. From the single bundle identity $D_{\mathcal{A}}\mathcal{F}=0$ we then derive the split into the Riemann and Yang–Mills Bianchi identities, $D_{\omega}R=0$ and $D_{A}F=0$, thereby establishing the unified Noether–II content. Variation with respect to $A$ yields the Yang–Mills equations $D_{A}(*_{h}F)=J_{\rm matter}$ Eq. (10), while variation with respect to $h$ yields Einstein’s equations sourced by the Yang–Mills and matter stress tensors Eq. (9). Therefore a single invariant action produces both sectors’ dynamics and their identities, exactly in Noether’s sense.
Within these principles there are no additional diffeomorphism- and gauge-natural, unitary second-order couplings in four dimensions beyond Einstein–Hilbert, Yang–Mills, and minimal matter terms as anything else is either purely topological, Euler, Pontryagin, $\theta$–terms or higher-dimension and therefore suppressed. Attempts to add an independent antisymmetric metric mode, extra massless higher-spin fields, or non-natural derivative couplings either break the symmetry structure, introduce ghosts or extra poles, or run afoul of anomaly constraints. Thus, in this construction, SM ++ GR is not an arbitrary choice but the unique possibility consistent with the Einstein–Noether program, and the Hermitian geometry makes that uniqueness geometrically explicit via $g_{[\mu\nu]}=F_{\mu\nu}$.

### v2

@@ 3 Real Slice Geometry
We will now explain real-slice geometry as a finer bundle to show how the real slice geometry encodes symmetries in HUFT. By a finer bundle over $M$ we mean an enlargement of a given bundle $q:P\to M$ by adjoining an extra fiber $E_{y}\to M$, encoding the imaginary directions $y$ from the complexification $M_{\mathbb{C}}$:
@@
so that fiberwise $\widehat{P}_{x}\cong P_{x}\times(E_{y})_{x}$ and $\pi_{1}$ simply forgets $y$. In our setting $P=P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$, so $\widehat{P}=(P_{\mathrm{Spin}}\times_{M}P_{G})\times_{M}E_{y}$. This is not a reduction in the structure group nor a refinement of an atlas, but it is an enrichment of the fiber used to keep track of the $y$-directions while all fields and dynamics are evaluated on the real slice $M$. We let $M$ be a smooth oriented, time-orientable spin $4$-manifold, the real slice. We have $M_{\mathbb{C}}$ as a complexification with coordinates $z^{\mu}=x^{\mu}+iy^{\mu}$ and projection $\pi:M_{\mathbb{C}}\to M$, $\pi(x,y)=x$. The fibers $Y_{x}:=\pi^{-1}(x)$ are real $4$-planes isomorphic to $\mathbb{R}^{4}$, defining a rank-$4$ real vector bundle $E_{y}\to M$ with total space canonically identified near $y=0$ with a neighborhood of $M\subset M_{\mathbb{C}}$.
We have $P_{Spin}\to M$ as the $Spin(1,3)$ frame bundle and $P_{G}\to M$ a principal $G$-bundle with $G=SU(3)\times SU(2)\times U(1)$. Now, we define the product bundle $P_{tot}:=P_{Spin}\times_{M}P_{G}$ with structure group $H:=Spin(1,3)\times G$. We refine this by adjoining the $y$-fiber:
@@
whose points encode local Lorentz frames, internal gauge frames, and the imaginary directions $y$ of the complexified manifold. Now, we choose an $H$-connection $\mathcal{A}=(\omega,A)$ on $P_{\mathrm{tot}}$ and an Ehresmann connection on $\pi:M_{\mathbb{C}}\to M$ that splits $TM_{\mathbb{C}}\simeq\mathsf{Hor}\oplus\mathsf{Ver}$ near $y=0$. Now, we write the unified curvature:
@@
and let $h\in\Gamma(\text{Sym}^{2}T^{*}M)$ be a Lorentzian metric, used for index operations and Hodge duals. On $M_{\mathbb{C}}$ we consider a Hermitian field:
@@
and identify on the real slice $B=F$ either by definition or via a holomorphic compatibility term enforcing BF=0B-F=0 in the equations of motion. The base metric $h$ transforms under $\text{Diff}(M)$ as its Levi–Civita and spin connection $\omega$ is the $Spin(1,3)$ part of $A$. The curvature $R$ is the gravitational field strength; all contractions, Hodge duals, and stress-energy are built using $h$. The internal symmetry $G=SU(3)\times SU(2)\times U(1)$ acts on $P_{G}$; the unified gauge potential $A=A^{(3)}\oplus A^{(2)}\oplus A^{(1)}$ and field strength $F=F^{(3)}\oplus F^{(2)}\oplus F^{(1)}$ live in $\Omega^{2}(M,\mathrm{ad}P_{G})$. On the real slice, the antisymmetric part of the Hermitian metric equals this curvature: $g_{[\mu\nu]}\equiv B_{\mu\nu}=F_{\mu\nu}$. For the Finer bundle with $y$ inside, the yy–fiber $E_{y}$ records the imaginary directions of $M_{\mathbb{C}}$. The Ehresmann splitting ties variations in $y$ to horizontal transport on $M$. In the holomorphic description, the mixed $(x,y)$-geometry packages the gauge sector into the antisymmetric two-form $B$ while preserving standard spacetime geometry in $h$. By Noether I and II, we obtain the single invariant action. A single $\text{Diff}(M)\times G$-invariant action:
@@
and by Noether’s second theorem, the unified Bianchi identities $D_{\omega}R=0$, $D_{A}F=0$, together with covariant current conservation $\nabla_{\mu}T^{\mu\nu}=0$, $D_{A}J=0$. The same symmetry explains the conservation laws. Writing $G=SU(3)\times SU(2)\times U(1)$ with Lie algebra $\mathfrak{g}=\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$ and Killing form $\kappa$, we decompose
@@
The diagram Figure 2 shows the complexification $\pi:M_{\mathbb{C}}\!\to M$, with the upper ribbon $M_{\mathbb{C}}$ and the lower ribbon is the real slice $M$. The point labelled $s$ is the image of the zero section $s_{0}:M\hookrightarrow M_{\mathbb{C}}$, $s_{0}(x)=(x,0)$; the lower point labelled $\pi$ is the basepoint $x\in M$, so $\pi\!\circ s_{0}=\mathrm{id}_{M}$. The curved line between them indicates the fiber $Y_{x}:=\pi^{-1}(x)$ over $x$, which is a real $4$-plane isomorphic to $\mathbb{R}^{4}$ and coordinatized by the imaginary directions $y^{\mu}$. We denote by $\mathbb{P}_{y}$ in the figure our $E_{y}$ in the text and the rank–$4$ real vector bundle $E_{y}\to M$ whose fiber at $x$ is $Y_{x}$; near $y=0$ its total space is canonically identified with a neighborhood of $M\subset M_{\mathbb{C}}$. This bookkeeping of the $y$-directions lets us speak of a refined finer bundle over $M$ that remembers the imaginary displacements while all fields are ultimately evaluated on the real slice.
Figure 2: The upper ribbon is $M_{\mathbb{C}}$ with coordinates $z^{\mu}=x^{\mu}+i\,y^{\mu}$, the lower ribbon is the real slice $M$. The projection $\pi:M_{\mathbb{C}}\!\to M$ sends $(x,y)\mapsto x$. The black dots mark the point $s_{0}(x)=(x,0)\in M_{\mathbb{C}}$ (label $s$) and its projection $x\in M$ (label $\pi$), so that $\pi\circ s_{0}=\mathrm{id}_{M}$. The label $\mathbb{P}_{y}$ denotes the rank-$4$ real vector bundle $E_{y}\to M$ of imaginary directions: for each $x\in M$, the fiber $Y_{x}=\pi^{-1}(x)\cong\mathbb{R}^{4}$ carries coordinates $y^{\mu}$.
@@ 4 Symmetry Completeness and Its Physical Equivalence
As a sanity check, we will show that we recover symmetry completeness and explore its physical equivalence. We have $M$ as an oriented, time-orientable spin 44–manifold and let $P_{\text{Spin}}\to M$ be the $\text{Spin}(1,3)$ frame bundle and $P_{G}\to M$ a principal $G$-bundle with $G=SU(3)\times SU(2)\times U(1)$. Set $P_{\text{tot}}=P_{\text{Spin}}\times_{M}P_{G}$ with structure group $H=\text{Spin}(1,3)\times G$. Now, let $h\in\Gamma(\text{Sym}^{2}T^{*}M)$ be a Lorentzian metric and $\mathcal{A}=(\omega,A)$ an $H$-connection with curvature $\mathcal{F}=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=(R,F)$, where $R$ is the Riemann curvature $2$-form and $F$ the Yang–Mills field strength. On the complexified bundle of $(0,2)$-tensors, define the Hermitian field as in Equation (70) and on the real slice impose the compatibility $B=F$ either as a definition or via a holomorphic penalty term enforcing BF=0B-F=0 in the equations of motion. We consider the single $\text{Diff}(M)\times G$-invariant action:
@@
where $\langle\cdot,\cdot\rangle_{h}$ is the fiberwise inner product induced by $h$ and the Killing form on $\mathfrak{g}$, and $\Psi$ denotes matter fields. Here, $\mathcal{A}=(\omega,A)$ with curvature $\mathcal{F}=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=(R,F)$, so $D_{\mathcal{A}}=(D_{\omega},D_{A})$ on $\mathfrak{spin}(1,3)\oplus\mathfrak{g}$.
{Theorem}
@@
    The Euler–Lagrange equations of $S$ are
@@
    (Classical equivalence of physics) The map
@@
    (Quantum equivalence, formal) If one includes a gauge-invariant penalty $\frac{\lambda}{2}\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$ and integrates out $B$, then, for $\lambda\to\infty$, the generating functional reduces to that of the standard Einstein–Yang–Mills and matter theory, so perturbative correlators and SS–matrix elements agree.
Under $H$-gauge transformations, $F\mapsto g^{-1}Fg$ and the Killing form is Ad-invariant, so $\mathrm{tr}(F\wedge*_{h}F)$ is gauge-invariant. Each term in (76) is a scalar density; hence, it is $\text{Diff}(M)$-invariant. For a connection $A$ on a principal $H$-bundle, $D_{A}F=0$ is the unified Bianchi identity, which splits as $D_{\omega}R=0$ and $D_{A}F=0$. Diffeomorphism invariance implies the contracted Bianchi identity $\nabla_{\mu}G^{\mu\nu}=0$ and hence $\nabla_{\mu}T^{\mu\nu}=0$ on shell $G$-invariance implies $D_{A}J=0$. Varying $A$ with $h$ fixed gives:
@@
Given any EYM++matter solution $(h,A,\Psi)$, define $(g,A,\Psi)$ by $g=h+iF$ and $\mathcal{A}=(\omega(h),A)$. Then, $(g,A,\Psi)$ solves the unified Euler–Lagrange system on the real slice because (77) hold, and $B=F$ by construction. Any unified real-slice solution $(g,A,\Psi)$ has $g=h+iB$ with $B=F$ from the BB–equation, and its $(h,A,\Psi)$ satisfies (77). Both constructions are natural with respect to $\text{Diff}(M)\times G$, so they descend to a bijection of quotient solution spaces. Classical observables built from $(h,A,\Psi)$, such as fluxes, charges, and stress-energy, coincide with those built from $(g,A,\Psi)$ restricted to the real slice. We add:
$S_{\lambda}[B,A;h]=\tfrac{\lambda}{2}\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h},$ (82)
the $B$-integral is Gaussian:
@@
HUFT is a unification of geometric data and variational origin rather than a claim of a new nonperturbative quantum-gravity principle. Classically, the real-slice theory is equivalent to Einstein–Yang–Mills and matter (Theorem 1), so HUFT makes the same classical gravitational predictions as GR coupled to the same gauge/matter content. Perturbatively, the formal quantum equivalence statement means that when the compatibility field $B$ is treated as auxiliary and integrated out, the generating functional reduces to that of standard EYM+matter up to an overall constant and identical gauge-fixing, so the perturbative spectrum and correlators are unchanged. For the UV completion and quantum-gravity aspect, if one seeks an ultraviolet-softened, BRST-compatible completion, the only allowed modifications consistent with the assumed local symmetries are covariant entire-function form factors $F(D^{2}/M_{\ast}^{2})$ in the kinetic operators as encoded in Theorem 2. In that sense, HUFT provides a symmetry-organized arena for discussing UV-finite nonlocal extensions of gravity+gauge theory, while remaining IR-equivalent to the local theory.
The consequences of this are that we have the same symmetries; $\text{Diff}(M)\times G$ invariance, Bianchi identities, and covariant current and stress-energy conservation are identical to GR++YM. The same observables, such as charges, fluxes, and classical predictions, match, so perturbative quantum correlators agree when $B=F$ is enforced. The unified packaging, gravity $R$ and gauge $F$, are components of one curvature $\mathcal{F}$ of one connection $\mathcal{A}$ on one bundle $P_{\mathrm{tot}}$.
{Theorem}
Let $(M,h)$ be a real Lorentzian $4$-manifold admitting a complexification $M_{\mathbb{C}}$. Let $P_{\mathrm{tot}}\to M$ be a principal $H$-bundle with $H=\mathrm{Spin}(1,3)\times G$ where $G$ is compact, reductive, and contains the SM gauge group. Consider fields $\big(h_{\mu\nu},\mathcal{A}=(\omega,A_{\mathrm{YM}}),\Psi,\Phi,g_{\mu\nu}\big)$ with
@@
Assume the dynamics are invariant under the automorphism group
@@
implemented off–shell by a nilpotent BRST differential $s$ (Noether II setting). The Lagrangian density is a diffeo and gauge-natural polynomial in the fields and their covariant derivatives, and may depend on the covariant d’Alembertian $D^{2}$ only through an entire functional calculus $F(D^{2}/M_{*}^{2})$ that commutes with $s$ and all covariant derivatives. In the limit $M_{*}\to\infty$, the quadratic operators for $(h,A,\Psi,\Phi)$ reduce to second order in derivatives, with indices raised and lowered solely by $h$. There is reflection positivity on the Euclidean slice as there are no extra propagating poles other than those of $(h,A,\Psi,\Phi)$. There are no additional massless higher-spin-gauge fields beyond spin–2 $h$ and spin–1 connections; these interactions are marginal or relevant in the IR. The local BRST anomaly cohomology $H^{1,4}(s\mid d)$ vanishes for the chosen matter representation, with no gauge or diffeo anomalies.
Then, modulo $s$-exact terms, total derivatives, and higher-dimension operators are suppressed by $M_{*}$, and the most general action is, after field redefinitions, equivalent to
@@
with the following consequences; BRST covariance implies the antisymmetric piece $B_{\mu\nu}$ transforms as a curvature two–form; hence, on the real slice,
$B\equiv F_{YM},\,\,\,\text{as elements of}\,\,\,\,\ \Omega^{2}(M,\mathrm{ad}\,P_{G}),$ (87)
where scalar contractions use $\kappa(\cdot,\cdot)$, so $g_{\mu\nu}=h_{\mu\nu}+iB_{\mu\nu}$ is a kinematic packaging and does not introduce an independent two-form gauge sector. Preserving the full local symmetry off–shell to all loops restricts UV softening to entire functions $F(D^{2}/M_{*}^{2})$ at covariant kinetic operators and consistently at vertices. This renders perturbation theory UV–finite while the IR ($M_{*}\to\infty$) recovers local EYM++matter. Noether II yields the unified Bianchi and Slavnov–Taylor identities, and
@@
with $T^{\mu\nu}$ the Belinfante tensor built from $h$. Up to $S_{\mathrm{top}}$, Euler, Pontryagin, $\theta$–terms, no further diffeo or gauge-natural, unitary, second-order couplings exist.
In this section, we specify the geometric data $P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$ with structure group $H=\mathrm{Spin}(1,3)\times G$, a single $H$-connection $\mathcal{A}=(\omega,A)$ with unified curvature $\mathcal{F}=(R,F)$, and a Lorentzian metric $h$. The master action Equation (19) is written and shown to be $\mathrm{Diff}(M)\times H$-invariant by construction. From the single bundle identity $D_{\mathcal{A}}\mathcal{F}=0$, we then derive the split into the Riemann and Yang–Mills Bianchi identities, $D_{\omega}R=0$ and $D_{A}F=0$, thereby establishing the unified Noether–II content. Variation with respect to $A$ yields the Yang–Mills equations $D_{A}(*_{h}F)=J_{\rm matter}$ Equation (22), while variation with respect to $h$ yields Einstein’s equations sourced by the Yang–Mills and matter stress tensors Equation (21). Therefore, a single invariant action produces both sectors’ dynamics and their identities, exactly in Noether’s sense.
Within these principles, there are no additional diffeomorphism- and gauge-natural, unitary second-order couplings in four dimensions beyond Einstein–Hilbert, Yang–Mills, and minimal matter terms, as anything else is either purely topological, Euler, Pontryagin, $\theta$-terms or higher-dimension and therefore suppressed. Attempts to add an independent antisymmetric metric mode, extra massless higher-spin fields, or non-natural derivative couplings either break the symmetry structure, introduce ghosts or extra poles, or run afoul of anomaly constraints. Thus, in this construction, SM ++ GR is not an arbitrary choice but the unique possibility consistent with the Einstein–Noether program, and the Hermitian geometry makes that uniqueness explicit by identifying the imaginary or antisymmetric sector of the Hermitian packaging tensor with an adjoint-valued two-form and enforcing real-slice compatibility:
$\mathrm{Im}(g_{\mu\nu})=\ell_{*}^{2}\,B_{\mu\nu},\qquad B\in\Omega^{2}(M,\mathrm{ad}\,P_{G}),$ (89)
The compatibility mechanism either imposed kinematically or derived dynamically enforces:
$B=F\quad\text{on the real slice }\,M,$ (90)
so that the Hermitian field $g=h+i\,\ell_{*}^{2}B$ is a packaging of $(h,F)$ rather than an additional propagating tensor sector. In particular, one should not write $g_{[\mu\nu]}=F_{\mu\nu}$ without specifying the projection to the imaginary/antisymmetric part and the scale $\ell_{*}$.